OpenStudy (anonymous):

HI THE FIRST ONE TO FINISH GETS MEDAL OR FAN YOU CHOOSE GL!! What is the surface area of a cube with side length 9 kilometers? two hundred sixteen square kilometers three hundred twenty-four square kilometers four hundred eighty-six square kilometers seven hundred twenty-nine square kilometers

OpenStudy (anonymous):

486 square kilometers

OpenStudy (anonymous):

9*9=81. (One Side) Cube has 6 sides, 81*6=486.

OpenStudy (anonymous):

congrats @andypie1010